OpenStudy (anonymous):

How do you work: Find the derivative of y= 2 cos x-3 sin x-7 (calculus question)

OpenStudy (anonymous):

use the chain rule i think.

OpenStudy (anonymous):

Thank you but i need a little bit more

OpenStudy (anonymous):

use the derivate

OpenStudy (anonymous):

@Andresfon12 . she is asking bout how to find it.

OpenStudy (mimi_x3):

Is it: \[2\cos(x)-3\sin(x)-7 \] or \[2\cos(x-3)\sin(x-7)\]?

OpenStudy (anonymous):

the 2nd one

OpenStudy (anonymous):

i guess u can use substitution for the x-3 and x-7